A jet bomber traveling 500 miles per hour arrived over its target at the same time as its fighter jet escort which left the same base 1 half hour after the bomber took off. How many hours did it take to reach the target if the fighter jet traveled 600 miles per hour? ( I couldn't write 1 half hour in fraction form so i just wrote it out.) ( I think you write it like this 1/2. But anyways can you help me please.)
Answer by mananth(12270)

let the target be at distance x
both arrive at same time
time taken by bomber = x/400
time taken by jet = x/500
jet took 1/2 hour less than bomber
x/400 -x/500=1/2
LCD=2000
multiply the equation by LCD
x*2000/400 - x*2000/500 = 1*2000/2
5x-4x=1000
x= 1000
speed of jet = 500 mph
Time taken by Jet = 1000/500= 2 hours
....
Time taken by Bomber = 1000/400 = 2.5 hours
m.ananth@hotmail.ca